A bobsled leaves a horizontal track with a speed of 30. m/s. It then moves onto a horizontal track lightly sprinkled with sawdust and comes to a stop in a distance of 150 m. Find the coefficient of friction between the runners of the sled and the sawdust-covered surface.
1
Expert's answer
2018-04-26T10:09:08-0400
From the conservation of energy:
(mv^2)/2=μmgd
μ=v^2/2gd
The coefficient of friction:
μ=〖30〗^2/(2(10)(150))=0.3.